You are a manager in a perfectly competitive market. The price in your market is $35. Your total cost curve is C(Q) = 10 + 2Q + .5Q2 and MC = 2 + Q.

a. What level of output should you produce in the short run?

b. What price should you charge in the short run?

c. Will you make any profits in the short run?

d. What will happen in the long run?
Answer #1

Giveny P= $35 C(Q) = 10 +2Q 2 to.5Q Mc = 2+Q output produce un Short run 35 = 2+Q 33 16-) Market price i.e. $35 charge in sho